# Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

### v4

@@ 3.2 Quality Problems in Responses (RQ2)
The average distribution of 12 problem types across four themes in responses to three repetitive experiments is illustrated in Fig. 5, where the same color represents types within the same theme, with darker shades indicating the themes. From the perspective of the four themes, most responses contain problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met. Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution. Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%). The typical types of problems within each theme are described in detail below.
Correctness: The responses reveal two types of incorrectness: Incorrect fact (6.25%) and Incorrect concept (10.13%). The responses with Incorrect fact include: incorrect line numbers of the code, function names not matching the source code provided, names not corresponding to the CWE-ID, and the described code content not matching the actual scenario. Incorrect concept indicates a misunderstanding of the concept of a certain CWE-ID or a specific security defect type. As shown in the example below (Example 1), the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
@@
Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%). Among these, Vague statement is the most prevalent one. It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing an overly general fix suggestion. An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
@@
Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses. Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appears in the responses generated by GPT-4. Such information includes the reasons for the absence of security defects in the current code and general suggestions on security measures and review processes. As illustrated in the example below (Example 3), the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence it is deemed to be Unnecessary information.
@@
Compliance: Three types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts. Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform a security check in its response or misinterpreted requirements in the prompts for other tasks. Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and fix suggestions of detected security defects requested in the prompt. The example below (Example 4) shows that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
@@
While Kabir et al. (2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours. Specifically, Kabir et al. (2024) requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues. In contrast, our study applied LLMs to security code review, revealing a much lower incidence of correctness issues at 14.81%. This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific application areas.
@@
We selected all items related to hallucination from the quality problem types in RQ1.1 and mapped them to the predefined set of hallucination subtypes. The details of mapping, corresponding explanations, and proportion results are shown in Table 7. Since a single response may exhibit multiple types of hallucination simultaneously, the sum of the proportions for each subtype exceeds the proportion of responses containing at least one hallucination. Overall, 20.19 % of the responses contain hallucinations. Among these, Factual Contradiction is the most frequent subtype, accounting for 10.13%, followed by Context Inconsistency and Factual Fabrication, which appear in 6.25% and 5.62% of responses, respectively. To alleviate these hallucinations, Retrieval-Augmented Generation (RAG) can be employed to incorporate external factual information (Tonmoy et al., 2024; Lewis et al., 2020), and post-training can also be applied to enhance the model’s capability to capture and retain contextual knowledge (Kumar et al., 2025).
@@
Factuality Hallucination Factual Contradiction Incorrect Concept In Incorrect Concept cases, the model’s output contradicts the authoritative definition of the corresponding CWE, thus categorized as Factual Contradiction. 10.13%
Factual Fabrication Groundless Speculation Groundless speculation refers to subjective assumptions made without contextual evidence or knowledge support. As it involves fabricated factual content, it is categorized as Factual Fabrication. 5.62%
Faithfulness Hallucination Instruction Inconsistency Detection Failure Detection Failure refers to cases where the model fails to perform the security checks or misinterprets the task, leading to deviations from the instruction. Therefore, it is categorized as Instruction Inconsistency. 0.15%
Context Inconsistency Incorrect Fact Incorrect Fact denotes cases where the line numbers, identifiers, or code content generated by the model do not align with the provided code context, thus categorized as Context Inconsistency. 6.25%
Logical Inconsistency Logical Contradiction Logical Contradiction refers to cases where the model’s reasoning about a security defect is internally contradictory, thus classified as Logical Inconsistency. 1.11%
Any Hallucination 20.19%
@@ 3.3.1 Model Stability
To assess the adequacy of the fit of our regression model, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991). This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022). A possible explanation is that although we have considered the non-determinism of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset. Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model. Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
@@
Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(>Chi)
Token 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
Community Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
FileType Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
AnnotationisSecurityRelated -1:0 0.97 2 9.95 0.0069**
1:0 2.69
AuthorExperience 0.87 1 4.89 0.0271*
Commit 1.01 1 3.06 0.0801
SecurityDefectType Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
Crash:Race Condition 0.86
Deadlock:Race Condition 0.92
DoS:Race Condition 0.87
Encryption:Race Condition 1.18
Improper Access:Race Condition 0.68
Integer Overflow:Race Condition 0.86
Resource Leak:Race Condition 1.41
Use After Free:Race Condition 0.37
Extra:Race Condition 1.43
Complexity 1.03 1 1.74 0.1868
AnnotationRatio 0.56 1 1.55 0.2126
AnnotationHasCode 1.11 1 1.09 0.2973
@@ 3.3.2 Explanatory Power of Each Factor
The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable. A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model. We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact. As shown in Table 8, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (<.0001). Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$<.0001) and ($\chi^{2}$=18.28, $p$<.0001). Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively. However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold. Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom. As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
@@ 3.3.3 Relationships between Factors and Responses
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 8). The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively. This suggests that the LLM is more likely to detect security defects in code files with fewer tokens, that contain functional logic, and were written by developers with less involvement in the project. Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt. In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations, the LLM is more likely to identify the target security defect. Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97. This suggests that annotations may disrupt the LLM and are less likely to detect the target security defect.

### v5

@@ 3.2 Quality Problems in Responses (RQ2)
[…]
For GPT-4, most responses contained problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involved Compliance-related problems. Furthermore, 27.08% of the responses showed problems with Understandability and problems related to Correctness were the least frequent, occurring in only 14.08% of the responses. In the responses of DeepSeek-R1, Conciseness problems were also the most prevalent, accounting for 84.29%. However, the frequencies of problems related to Compliance and Understandability were significantly lower than those of GPT-4, at 4.51% and 9.54%, respectively. Moreover, DeepSeek-R1 was much more prone to Correctness problems than GPT-4, with a rate of 33.66%. The typical types of problems within each theme are described in detail below.
Correctness: The responses revealed two types of incorrectness: Incorrect fact (GPT-4: 6.25%, DeepSeek-R1: 31.81%) and Incorrect concept (GPT-4: 10.13%, DeepSeek-R1: 2.39%). The responses with Incorrect fact include: incorrect line numbers of the code, function names not matching the source code provided, names not corresponding to the CWE-ID, and the described code content not matching the actual scenario. Incorrect concept indicates a misunderstanding of the concept of a certain CWE-ID or a specific security defect type. As shown in the example below (Example 1), the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem. Compared with GPT-4, DeepSeek-R1 exhibited a lower frequency of Incorrect concept problems, but its Incorrect fact rate was significantly higher. We observed that during security code reviews, GPT-4 tended to provide the range of code lines where defects were located, while DeepSeek-R1 often identified specific code lines and their corresponding code, making it more prone to errors in details. This may contribute to the higher Incorrect fact rate of DeepSeek-R1 compared to GPT-4. When applying LLMs to real-world security code reviews, we expect models to provide both specific and accurate factual information. Therefore, how to enhance the capability of LLMs to retain code details with precision requires further exploration.
@@
Understandability: Problems that reduce understandability of responses include Vague statement (GPT-4: 21.38%, DeepSeek-R1: 3.58%), Incoherent expression (GPT-4: 1.19%, DeepSeek-R1: 0.53%), Logical contradiction (GPT-4: 1.11%, DeepSeek-R1: 1.13%), and Groundless speculation (GPT-4: 5.62%, DeepSeek-R1: 4.11%). For GPT-4, Vague statement is the most prevalent one. It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing an overly general fix suggestion. An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response. While for DeepSeek-R1, the frequency of Vague statement was considerably lower than that of GPT-4. We speculated that its reasoning procedure under the ‘think’ mode somewhat contributed to generating specific and detailed responses.
@@
Conciseness: In all responses, some of them contain Redundant content (GPT-4: 1.14%, DeepSeek-R1: 39.76%), some include information Irrelevant to the task of security defect detection (GPT-4: 15.12%, DeepSeek-R1: 6.03%), and some contain Unnecessary elements (GPT-4: 86.38%, DeepSeek-R1: 63.29%), all of which are problems related to the conciseness of the response. We found that DeepSeek-R1 was more likely than GPT-4 to repeatedly mention the same information in its output, whereas GPT-4 produced content that is unrelated to the target task more frequently. Including Unnecessary content, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, is the most prevalent problem for both GPT-4 and DeepSeek-R1. Such information includes the reasons for the absence of security defects in the current code and general suggestions on security measures and review processes. As illustrated in the example below (Example 3), the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence it is deemed to be Unnecessary information.
@@
Compliance: Four types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts. Among them, Detection failure was observed in very few responses (GPT-4: 0.15%, DeepSeek-R1: 0.13%), where LLM explicitly claimed that it could not perform a security check in its response or was stuck in an infinite output loop and failed to complete the security defect detection task. Task misunderstanding, which is similarly rare (GPT-4: 0.00%, DeepSeek-R1: 0.66%), indicates that LLM misinterprets the task details in the prompt. For instance, when using DeepSeek-R1 with $\textbf{P}_{cot-guardrail}$, if the commit message mentioned security defects addressed by the current commit, LLM may mistakenly interpret the task of security code review as explaining the security defects mentioned in the commit message. Non-compliant output was present in 20.43% of GPT-4 responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme for GPT-4, referring to cases in which the responses lack the line numbers or fix suggestions of detected security defects requested in the prompt. Although GPT-4 showed relatively high frequencies for these two problems, DeepSeek-R1 exhibited much lower rates for them, at 1.86% and 1.99%, respectively. Given this, we consider DeepSeek-R1 to exhibit stronger instruction-following capabilities than GPT-4 in security code review. The example below (Example 4) shows that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
@@
While kabir2023answers also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours. Specifically, kabir2023answers requested ChatGPT to answer programming questions on Stack Overflow and found that ChatGPT makes fewer factual errors compared to conceptual errors. This aligns with our findings for GPT-4. However, unlike their study, which identified correctness problems in over half (52%) of the responses, our study applied GPT-4 to security code review, revealing a much lower incidence of correctness issues at 14.81%. When applying DeepSeek-R1, the frequency of correctness problems increased to 33.66%. This suggests that the distribution of quality problems in LLM responses may vary significantly across different application domains and model architectures, necessitating distinct optimization directions and strategies in practical application.
@@
We selected all items related to hallucination from the quality problem types in RQ1.1 and mapped them to the predefined set of hallucination subtypes. The details of mapping, corresponding explanations, and proportion results are shown in Table 7. Since a single response may exhibit multiple types of hallucination simultaneously, the sum of the proportions for each subtype exceeds the proportion of responses containing at least one hallucination. Overall, 20.19% of GPT-4 responses contained hallucinations, whereas DeepSeek-R1 exhibited a higher rate of 36.45%. The former encountered more Factuality Hallucination, while the latter was predominantly affected by Faithfulness Hallucination. Specifically, for GPT-4, Factual Contradiction was the most frequent subtype, accounting for 10.13%, followed by Context Inconsistency (6.25%) and Factual Fabrication (5.62%). In contrast, for DeepSeek-R1, Context Inconsistency was the most prevalent subtype (31.81%), with Factual Fabrication and Factual Contradiction following at 4.11% and 2.39%, respectively. To alleviate these hallucinations, Retrieval-Augmented Generation (RAG) can be employed to incorporate external factual information (tonmoy2024comprehensive; lewis2020retrieval), and post-training can also be applied to enhance the model’s capability to capture and retain contextual knowledge (kumar2025llm).
@@
GPT-4 DS-R1
Factuality Hallucination Factual Contradiction Incorrect Concept In Incorrect Concept cases, the model’s output contradicts the authoritative definition of the corresponding CWE, thus categorized as Factual Contradiction. 10.13% 2.39%
Factual Fabrication Groundless Speculation Groundless speculation refers to subjective assumptions made without contextual evidence or knowledge support. As it involves fabricated factual content, it is categorized as Factual Fabrication. 5.62% 4.11%
Faithfulness Hallucination Instruction Inconsistency Detection Failure, Task Misunderstanding Detection Failure refers to cases where the model fails to perform the security checks or misinterprets the task, leading to deviations from the instruction. Therefore, it is categorized as Instruction Inconsistency. 0.15% 0.80%
Context Inconsistency Incorrect Fact Incorrect Fact denotes cases where the line numbers, identifiers, or code content generated by the model do not align with the provided code context, thus categorized as Context Inconsistency. 6.25% 31.81%
Logical Inconsistency Logical Contradiction Logical Contradiction refers to cases where the model’s reasoning about a security defect is internally contradictory, thus classified as Logical Inconsistency. 1.11% 1.13%
Any Hallucination 20.19% 36.45%
@@ 3.3.1 Model Stability
[…]
@@
Factor (:Reference) GPT-4 DeepSeek-R1
OR D.F. $\chi^{2}$ Pr(>Chi) OR D.F. $\chi^{2}$ Pr(>Chi)
Token 0.63 1 54.20 $0001^{***}$ 0.44 3 9.81 0.0017**
Position 0.89 1 0.2685 0.6042 1.15 1 0.47 0.4909
Community Qt:Openstack 1.54 1 8.14 0.0043** 1.41 1 6.03 0.0141*
FileType Source:Auxiliary 2.81 1 19.04 $0001^{***}$ 0.57 1 5.57 0.0183*
AnnotationisSecurityRelated -1:0 1.04 2 15.46 0.0004*** 1.28 2 14.02 0.0009***
1:0 2.91 2.57
AuthorExperience 0.93 1 1.29 0.2569 0.96 1 0.65 0.4203
Commit 1.01 1 4.82 0.0281* 1.00 1 0.01 0.7379
SecurityDefectType Buffer Overflow:Race Condition 0.84 10 11.29 0.3353 1.25 10 28.12 0.0017**
Crash:Race Condition 0.94 1.01
Deadlock:Race Condition 0.83 0.93
DoS:Race Condition 0.98 1.99
Encryption:Race Condition 0.83 4.43
Improper Access:Race Condition 0.80 2.16
Integer Overflow:Race Condition 0.85 1.16
Resource Leak:Race Condition 1.49 1.14
Use After Free:Race Condition 0.41 0.84
Extra:Race Condition 0.74 1.99
Complexity 1.03 1 2.60 0.1071 1.08 1 15.51 $0001^{***}$
AnnotationRatio 0.60 1 1.30 0.2538 1.16 1 0.11 0.7379
AnnotationHasCode 0.83 1 1.03 0.3094 1.46 1 4.37 0.0364*
@@ 3.3.2 Explanatory Power of Each Factor
The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable. A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model. We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact.
As shown in Table 8, for GPT-4, ‘Token’ wielded the highest predictive power on the fitted model, with the highest $\chi^{2}$ (54.20) and the lowest p-value (<.0001). Trailing closely behind were ‘FileType’ and ‘AnnotationisSecurityRelated’ factors, with correspondingly significant contributions ($\chi^{2}$=19.04, pp<.0001) and ($\chi^{2}$=15.46, pp=0.0004). Adding to this, ‘Community’ and ‘Commit’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 8.14 and 4.82, respectively. For DeepSeek-R1, ‘Complexity’ and ‘AnnotationisSecurityRelated’ demonstrated the strongest explanatory power over the response variable, with $\chi^{2}$ of 15.51 and 14.02, and p-values both less than 0.001. Following closely were ‘SecurityDefectType’ and ‘Token’, with p-values of 0.0017 (less than 0.01), indicating that they also have a statistically significant impact on the model’s performance in security code review. Although the $\chi^{2}$ for ‘Community’ (6.03), ‘FileType’ (5.57) and ‘AnnotationHasCode’ (4.37) were relatively smaller, they still contributed predictive value in the fitted model. Factors not analyzed here were deemed not to largely influence the performance of GPT-4 and DeepSeek-R1 in detecting security defects in code reviews, as their p-values exceeding the 0.05 threshold and $\chi^{2}$ values were much lower than those discussed factors. Notably, ‘SecurityDefectType’ had a relatively high $\chi^{2}$ value (11.29 for GPT-4 and 28.12 for DeepSeek-R1), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom. As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
@@ 3.3.3 Relationships between Factors and Responses
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (paul2021security; thongtanunam2017review; setia2016methodology). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (harrell2017package) to calculate the OR for each factor (see Table 8).
As presented in Table  8, for both DeepSeek-R1 and GPT-4, the OR for the ‘token’ factor was less than 1, indicating that the models are more likely to detect security defects in code files with fewer tokens. The OR values for ‘AnnotationisSecurityRelated’ and ‘community’ were greater than 1, suggesting that the two models are more likely to detect security defects in code files that include security-related annotations and originate from the Qt community, rather than from OpenStack. However, regarding ‘FileType’, GPT-4 and DeepSeek-R1 showed opposing results. GPT-4 had an OR of 2.81 (>1>1), indicating a higher likelihood of detecting security defects in source code with functional logic, while DeepSeek-R1 had an OR of 0.57 (<1<1), suggesting better performance on header files utilized for declaration. Furthermore, DeepSeek-R1 exhibited varying detection capabilities for different types of security defects. According to the OR values of ‘SecurityDefectType’, DeepSeek-R1 demonstrated a greater tendency to detect various other security defect types compared to Race Condition, including but not limited to Encryption, Buffer overflow, Integer Overflow, and Resource leak. In addition, DeepSeek-R1 displayed stronger capabilities in security defect detection for files with more code complexity, as evidenced by its OR for ‘complexity’ being greater than 1 (1.08). This contrasts with the consensus in previous works that as complexity of software grows, security defects are becoming more challenging to detect sheng2024lprotector. A possible explanation is that we utilized McCabe’s Cyclomatic Complexity mccabe1976complexity to measure this factor, which is not always positively correlated with the actual difficulty of detection. Code with higher cyclomatic complexity may have a clearer code structure and retain more contextual information within the single code file, which could better support DeepSeek-R1’s reasoning procedure during its think mode. On the other hand, the analysis of security defects in code with lower cyclomatic complexity may require more cross-file code context and project-specific knowledge, leading to DeepSeek-R1 struggles to determine whether security defects exist.
It is also worth noting that while previous studies have explored the relationship between the position of security defects and the performance of LLMs, our findings differ. According to the study by Sovrano_2025, the ‘position’ of security defect was determined by counting the characters preceding the security defect and they concluded that security defects located at the end of the code file are often harder to detect. However, when the file itself contains a large number of characters, even if the defect is located in the relatively middle or beginning part of the file, its ‘position’ can still be quite high. Therefore, the logistic regression analysis in Sovrano_2025 was to some extent confounded by the effect of character count (or token count). Our study, by measuring ‘position’ as the ratio of tokens preceding the security defect relative to the total token count, reduced the correlation between ‘Position’ and ‘Token’. It can be seen from Table 8 that in both of the two fitted models, the p-values of ‘Position’ were far greater than 0.05, indicating that the position of security defect has no significant effect on the performance of either GPT-4 or DeepSeek-R1. While the OR of ‘Position’ (0.89) in GPT-4 was less than 1, consistent with  Sovrano_2025, DeepSeek-R1 showed an OR of 1.15, which is greater than 1. Therefore, the impact of the position of security defect in code files on the detection performance of latest reasoning models warrants further analysis and validation.